A person traveling 4 hrs by plane and 25 hrs by ship covers 1580 miles. If the speed of plane had been one-half of the actual speed and the speed of the ship had been one-fourth greater, the person would have traveled only 1315 miles in the same length of time. Find the speeds of the plane and the ship.
Thanks A Bunch!

Answer by venugopalramana(3286)

A person traveling 4 hrs by plane and 25 hrs by ship covers 1580 miles.
LET THE SPEED OF PLANE=P MPH
LET THE SPEED OF SHIP =S MPH
DISTANCE TRAVELLED BY PLANE=4P
DIST.TRVLD.BY SHIP=25S
TOTAL DIST.TRVLD.=4P+25S=1580......................I
If the speed of plane had been one-half of the actual speed and the speed of the ship had been one-fourth greater, the person would have traveled only 1315 miles
NEW SPEED OF PLANE =P/2
NEW SPEED OF SHIP..1/4 TH.GREATER = S+S/4=5S/4
DIST.TRVLD.=4P/2+25*5S/4=1315
8P+125S=5260.............................II
EQN.II - EQN.I*2
125S-50S=5260-3160=2100
75S=2100
S=2100/75=28 MPH
P=(1580-25*28)/4=220
in the same length of time. Find the speeds of the plane and
the ship.
Thanks A Bunch!